OpenStudy (awesomedude2078):

Calculate the upper and lower limit for a 95% confidence interval about the mean. A family wants to reduce its expenditures for personal items like gifts, newspapers, magazines and so forth. A sample of 49 months of receipts yields a mean of $220.00 with a standard deviation of $30.00. They decide to calculate a 95% confidence interval about this mean. Standard error = (standard deviation)/(square root of sample size) Upper limit (dollars and cents) _________ Lower limit (dollars and cents) _________

OpenStudy (t-dawg02):

Upper limit (dollars and cents) $228.40 Lower limit (dollars and cents) $211.60

OpenStudy (t-dawg02):

Need an explanation?

OpenStudy (awesomedude2078):

yeah i don't understand how to do this

OpenStudy (t-dawg02):

Okay

OpenStudy (t-dawg02):

Standard error = (standard deviation)/(square root of sample size) = 30/√49 = 30/7 1.96 * standard error = 1.96 * 30/7 = 8.4 Upper limit = mean + 1.96SE = 220 + 8.4 Lower limit = mean - 1.96SE = 220 - 8.4

OpenStudy (awesomedude2078):

ok i see now, thank you

OpenStudy (t-dawg02):

Anytime :)